pi/4 ∫ 0 sec x dx

pi/40 sec x dx

Grade:12

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

Assume sinx = t,
cosx dx = dt,
so, integration becomes\int_{0}^{1/\sqrt{2}}1 / (1-t2) dt

Solove this integral by writing (1-t2) = (1-t) (1+t)

thnaks.